OpenStudy (bossimbacon):

What value of w solves the equation? w^-3= 1/64 will give fan and medal

OpenStudy (bossimbacon):

@melissa780 could you please help

OpenStudy (anonymous):

w^3 = 1/64 w = 1/4 because 1/4 ^3 = 1/64 1/64 = 1/64 Is 1/4 one of your options?

OpenStudy (bossimbacon):

it just shows a box and tells me to put the value in

OpenStudy (bossimbacon):

so whatever you think is right i will put it in :D

OpenStudy (bossimbacon):

oh whoops i put the equation wrong sorry w^-3 = 1/64

OpenStudy (anonymous):

Are you sure it's that?

OpenStudy (bossimbacon):

yes

OpenStudy (bossimbacon):

i have to go soon:( can you tell me plz

OpenStudy (anonymous):

Sorry, don't know that one.

OpenStudy (bossimbacon):

:( sucks well i have to go thanks for trying

OpenStudy (anonymous):

You're welcome, sorry.

OpenStudy (bossimbacon):

its fine!